OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = -x2 + 8x + 2 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 1 to x = 4 represent?

OpenStudy (anonymous):

@dan815 @texaschic101 @geerky42 @sammixboo

OpenStudy (anonymous):

@DarkMoonZ @Donblue22 @DaWizjr

OpenStudy (anonymous):

I know it's c or d

OpenStudy (anonymous):

I think c

OpenStudy (anonymous):

@clynnew

OpenStudy (anonymous):

@dan815 please help

OpenStudy (welshfella):

where are the options?

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

The water travels an average distance of 3 feet from 1 second to 4 seconds. The water travels an average distance of 2 feet from 1 second to 4 seconds. The water rises up with an average speed of 3 feet per second from 1 second to 4 seconds. The water rises up with an average speed of 2 feet per second from 1 second to 4 seconds.

OpenStudy (anonymous):

r u still there?

OpenStudy (welshfella):

the graph will be a parabola which opens downwards and will have a maximum value when x = 1 f(x) = -1 + 8 + 2 when x = 4 f(x) = -16 + 32 + 2 these are the heights at these times

OpenStudy (welshfella):

so calculate the distance travelled upwards for this period and divide by the time (3 seconds)

OpenStudy (anonymous):

wait why -1 and -16? Isn't -4 times -4 16?

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@Phebe @sammixboo @bohotness @geerky42